# On the semi-additivity of the $1/2$-symmetric caloric capacity in the plane

### v3

@@ 3.3. An equivalent variational capacity
$\gamma_{\text{{sy}},\psi_{\tau},+}(E)$ $\approx\gamma_{\text{{sy}},\psi_{\tau},+}^{\prime}(E)\approx\gamma_{\text{{sy}},\psi_{\tau},+}^{\star}(E)\approx\gamma_{\text{{sy}},\psi_{\tau},2}(E)\approx\gamma_{\text{{sy}},\psi_{\tau},\text{{op}}}(E)$
as well as
$\limsup_{\tau\to 0}\gamma_{\text{{sy}},\psi_{\tau},+}(E)\lesssim\gamma_{\text{{sy}},+}(E)\lesssim\liminf_{\tau\to 0}\gamma_{\text{{sy}},\psi_{\tau},+}(E).$
The latter relations for the capacities depending on $\psi_{\tau}$ also hold changing each $\gamma_{\text{{sy}}}$ for $\widetilde{\gamma}$.
@@ 4. The construction of the cubes
Let $H$ be a positive Borel measure supported on $E$ such that for any $\lambda_{0}>0$ and any $\lambda\in[0,\lambda_{0}]$, $\mu_{\lambda}:=\mu_{0}+\lambda H\in\Sigma(E)$. Then,
@@
$\geq\frac{\mu_{0}(E)^{2}+2\lambda\mu_{0}(E)H(E)+\lambda^{2}H(E)^{2}}{\mu_{0}(E)+\lambda H(E)+\int_{E}(\pazocal{S}\mu_{0})^{2}\mathop{}\!\mathrm{d}\mu_{0}+\lambda\big{[}\cdots\big{]}+\lambda^{2}\big{[}\cdots\big{]}+\lambda^{3}\big{[}\cdots\big{]}},$
@@
Assume that $\lambda>0$ and divide both sides of the previous inequality by this factor and make $\lambda\to 0$ to obtain the desired estimate. Let us notice that the kernel associated to $\pazocal{S}$ is continuous, meaning that the terms in []delimited-[][$\cdots$][ ⋯ ] and $\int_{E}(\pazocal{S}H)^{2}\mathop{}\!\mathrm{d}H$ are finite, since $H(E)<\infty$ due to the fact that $\mu_{\lambda}\in\Sigma(E)$ for any $\lambda\in[0,\lambda_{0}]$ with $\lambda_{0}>0$. ∎
@@
We will now prove that for all sufficiently small $\lambda>0$, $\mu_{0}+\lambda\pazocal{L}^{n+1}|_{G}$ belongs to $\Sigma(E)$, i.e.
@@
where for the last inequality we have applied the reasoning of the second case. Hence, in general, we deduce that for $\lambda>0$ small enough, $\mu_{\lambda}\in\Sigma(E)$.
@@ 4.1. The construction
since $P_{\text{{sy}}}$ and the measure $\mu$ are nonnegative. Therefore, $\pazocal{S}_{\mu}$ becomes a C-Z convolution operator (with continuous kernel), so that it satisfies the weak estimate
@@
Therefore, using the C-Z estimates for $P_{\text{{sy}}}\psi_{\tau_{0}}$, the nonnegativity of the latter kernel and relations (4.3) we have
@@ 1. $\text{{{A}}}_{\boldsymbol{1}}$:
    $E$ is contained in the unit ball and consists of a finite union of dyadic cubes belonging to a dyadic grid in $\mathbb{R}^{n+1}$ ((((with sides parallel to the coordinate axes))\,)), all with the same size and with disjoint interiors.
Let us verify that if we deduce the comparability between $\widetilde{\gamma}$ and $\widetilde{\gamma}_{+}$ for $E$ satisfying $\text{{{A}}}_{\boldsymbol{2}}$, we obtain the same result for a general $E$. So fix $E\subset\mathbb{R}^{n+1}$ any compact set and let $\overline{\pazocal{U}_{\delta}(E)}$ be the closed $\delta$-neighbourhood of $E$. Consider a grid of dyadic cubes in $\mathbb{R}^{n+1}$ with sides parallel to the axes and diameter smaller than $\delta/2$. Name $E_{0}$ the collection of dyadic cubes that intersect $E$ and notice that $E\subset E_{0}\subset\overline{\pazocal{U}_{\delta}(E)}$. Now we would have
$\widetilde{\gamma}(E)\leq\widetilde{\gamma}(E_{0})\lesssim\widetilde{\gamma}_{+}(E_{0})\leq\widetilde{\gamma}_{+}(\overline{\pazocal{U}_{\delta}(E)})\leq\gamma_{+}(\overline{\pazocal{U}_{\delta}(E)}).$
@@
Finally, we present the final additional assumption that motivates the title of this section:
@@
The aim of this section is to prove the comparability between $\widetilde{\gamma}$ and $\widetilde{\gamma}_{+}$ under assumption $\text{{{A}}}_{\boldsymbol{3}}$. Later on, we will get rid of these hypotheses in some particular cases, for which we will obtain the desired comparability in full generality.
@@ 5.1. Basic definitions and properties
$\bigg{\rvert}\int_{Q}\varphi(\overline{x})\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})\bigg{\rvert}\leq\bigg{\rvert}\int_{Q}\partial_{t}(P\ast\varphi)(\overline{x})\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})\bigg{\rvert}+\bigg{\rvert}\int_{Q}(\Delta_{x})^{1/2}(P\ast\varphi)(\overline{x})\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})\bigg{\rvert}$
@@
$\bigg{\rvert}\int_{Q}\partial_{t}(P\ast\varphi)(\overline{x})\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})\bigg{\rvert}$ $=\bigg{\rvert}\int_{Q_{1}}\bigg{(}\int_{a}^{b}\partial_{t}(P\ast\varphi)(x,t)\mathop{}\!\mathrm{d}t\bigg{)}\mathop{}\!\mathrm{d}\pazocal{L}^{n}(x)\bigg{\rvert}$
$=\bigg{\rvert}\int_{Q_{1}}\Big{(}P\ast\varphi(x,b)-P\ast\varphi(x,a)\Big{)}\mathop{}\!\mathrm{d}\pazocal{L}^{n}(x)\bigg{\rvert}$
$\lesssim\|P\ast\varphi\|_{\infty}\pazocal{L}^{n}(Q_{1})\leq\ell(Q)^{n}.$
@@
$(\Delta_{x})^{1/2}\approx\sum_{j=1}^{n}\partial_{j}R_{j}$
[…]
@@ 5.2. The exceptional sets $\boldsymbol{H_{\pazocal{D}}}$ and $\boldsymbol{T_{\pazocal{D}}}$
Pick a subfamily of disjoint maximal cubes $\{Q_{k}\}_{k}$ from $\pazocal{D}_{H}$ so that we may define the exceptional set $H{D}$ as follows
@@
We define the exceptional set $T{D}$ in a similar manner but imposing a proper accretivity condition on the Radon-Nikodym derivative $b$. Consider the family $\pazocal{D}_{T}\subset\pazocal{D}$ of cubes satisfying
@@
5.3. Verifying the hypotheses of a $Tb\,$-theorem
The goal of this subsection is to prove the lemma below, analogous to [To3, Lemma 6.8 (Main Lemma)]. We remark that in order to apply the $Tb$-theorem found in [NTVo1] we will still need to check an additional weak boundedness property for a particular suppressed kernel, due to the lack of anti-symmetry of our kernel $P$. Following the proof of the aforementioned reference or that of [To3, Theorem 5.1], one notices that such condition needs not to be verified, since the Cauchy kernel satisfies being anti-symmetric.
@@ 5.3. Verifying the hypotheses of a $Tb\,$-theorem
    If $\mu(B_{r}(\overline{x}))>Lr^{n}$ (((\,(for some big constant $L)$, then $B_{r}(\overline{x})\subset H$. In particular, for any $\overline{x}\in F\setminus{H}$ and $r>0$, one has $\mu(B_{r}(\overline{x}))\leq Lr^{n}$.
@@
Let us check that many of the hypothesis in Lemma 5.3 are verified for our particular choice of $\mu$ and $\nu$. Properties 1, 2 and 4 are proved in Lemma 5.1, and 3 follows from the third property of the same lemma, and by assumption $\text{{{A}}}_{\boldsymbol{3}}$ combined with $r_{i}^{n}\approx\widetilde{\gamma}(2Q_{i}\cap E)$. Moreover, taking $H\subset F$ to be as in (5.6), properties 6 and 7 are also guaranteed by definition. So it remains to check hypothesis 5. To do so, let us begin by presenting a series of auxiliary results.
@@
To prove 2, assume that $\|P\ast T\|_{\infty}\leq 1$ and $\|P^{\ast}\ast T\|_{\infty}\leq 1$. Consider $\psi$ as before and define the auxiliary function
@@
that is such that $\|\nabla\eta\|_{\infty}\leq\ell(2Q)^{-1}$. By Remark 2.1, we shall apply [MPr, Theorem 3.1] to deduce $\|P\ast\eta T\|_{\infty}\lesssim 1$ and $\|P^{\ast}\ast\eta T\|_{\infty}\lesssim 1$. Hence $|\langle\eta T,1\rangle|\lesssim\widetilde{\gamma}(E)$, that by definition means
@@
$|P\ast\widetilde{\varphi_{i}T_{0}}|(\overline{x})\leq\frac{|\langle T_{0},\varphi_{i}\rangle|}{\pazocal{L}^{n+1}(B_{i})}\int_{B_{i}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}\lesssim\frac{1}{r_{i}}\int_{B_{i}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}},$
@@
$\int_{B_{i}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}$ $=\int_{B_{i}\cap\{|\overline{x}-\overline{y}|\geq r_{i}\}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}+\int_{B_{i}\cap\{|\overline{x}-\overline{y}|<r_{i}\}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}$
$\leq\frac{\pazocal{L}^{n+1}(B_{i})}{r_{i}^{n}}+\int_{B_{4r_{i}}(\overline{x})}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}\lesssim r_{i}.$
@@
(5.10) $\hskip 56.9055pt+\frac{|\langle\varphi_{i}T_{0},1\rangle|}{\pazocal{L}^{n+1}(B_{i})}\int_{B_{i}}\big{\rvert}P(\overline{z}-\overline{y})-P(\overline{z}-c_{B_{i}})\big{\rvert}\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})$
@@
$\frac{|\langle\varphi_{i}T_{0},1\rangle|}{\pazocal{L}^{n+1}(B_{i})}\int_{B_{i}}\big{\rvert}P(\overline{z}-\overline{y})-$ $P(\overline{z}-c_{B_{i}})\big{\rvert}\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})$
$\lesssim\frac{1}{r_{i}}\frac{1}{|\overline{z}-c_{B_{i}}|^{n+1}}\int_{B_{i}}|\overline{y}-c_{B_{i}}|\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})$
@@
$R_{\varepsilon}(\overline{x}_{0})=P(\overline{x}_{0})\quad\text{ and }\quad R_{\varepsilon}^{\ast}(\overline{x}_{0})=P^{\ast}(\overline{x}_{0}).$
@@
$|R_{\varepsilon}(\overline{x}_{0})|\lesssim\frac{1}{\varepsilon^{n}}\quad\text{ and }\quad|R_{\varepsilon}^{\ast}(\overline{x}_{0})|\lesssim\frac{1}{\varepsilon^{n}}.$
@@
$|R_{\varepsilon}(\overline{x}_{0})|\leq\frac{1}{\varepsilon^{n+1}}\int_{B_{\varepsilon}(0)}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})}{|\overline{x}_{0}-\overline{x}|^{n}}\leq\frac{1}{\varepsilon^{n+1}}\int_{2B_{\varepsilon}(\overline{x}_{0})}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})}{|\overline{x}_{0}-\overline{x}|^{n}}\lesssim\frac{1}{\varepsilon^{n}},$
@@
$|R_{\varepsilon}\ast\alpha_{i}(\overline{z})|\leq\int_{\text{supp}(\Psi_{\varepsilon}\ast\alpha_{i})}P(\overline{z}-\overline{y})\big{\rvert}\Psi_{\varepsilon}\ast\alpha_{i}(\overline{y})\big{\rvert}\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y}).$
@@
$|R_{\varepsilon}\ast\alpha_{i}(\overline{z})|$ $\leq\|\Psi_{\varepsilon}\ast\alpha_{i}\|_{\infty}\int_{\pazocal{U}_{\varepsilon}(2Q_{i})}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{z}-\overline{y}|^{n}}\lesssim(\ell(Q_{i})+\varepsilon)\cdot\|\Psi_{\varepsilon}\ast\alpha_{i}\|_{\infty}$
@@
$=\sum_{i=1}^{N}\bigg{(}\int_{4Q_{i}}R_{\ast}\ast\big{(}\widetilde{\varphi_{i}T_{0}}-\varphi_{i}T_{0}\big{)}(\overline{y})\mathop{}\!\mathrm{d}\mu(\overline{y})+\int_{F\setminus{(4Q_{i}\cup H)}}R_{\ast}\ast\big{(}\widetilde{\varphi_{i}T_{0}}-\varphi_{i}T_{0}\big{)}(\overline{y})\mathop{}\!\mathrm{d}\mu(\overline{y})\bigg{)}$
$=:\sum_{i=1}^{N}I_{i,1}+I_{i,2},$
@@
The proof of 5 concludes that of Lemma 5.3. In order to apply the $Tb$-theorem of [NTVo1] or [To3, Theorem 5.1] we still need to check an additional weak boundedness property for our $n$-dimensional C-Z kernel, which fails being anti-symmetric.
@@
$\pazocal{S}_{1}^{\prime}$ :=$:=\big{\{}\overline{x}\in F\,:\,\pazocal{P}_{\ast}\nu(\overline{x})>\alpha\big{\}},$
$\pazocal{S}_{2}^{\prime}$ :=$:=\big{\{}\overline{x}\in F\,:\,\pazocal{P}^{\ast}_{\nu,\ast}1(\overline{x})>\alpha\big{\}},$
@@
$|\displaystyle\big{\rvert}\pazocal{P}_{\varepsilon_{0}(\overline{x})}\nu(\overline{x})-\pazocal{P}_{\varepsilon_{0}(\overline{x})}\nu(\overline{y})\big{\rvert}$$\leq\big{\rvert}\pazocal{P}_{\varepsilon_{0}(\overline{x})}\big{(}\chi_{B_{2\varepsilon_{0}(\overline{x})}(\overline{y})}\nu\big{)}\overline{x})\big{\rvert}+\big{\rvert}\pazocal{P}_{\varepsilon_{0}(\overline{x})}\big{(}\chi_{B_{2\varepsilon_{0}(\overline{x})}(\overline{y})}\nu\big{)}(\overline{y})\big{\rvert}$
@@
Regarding the third term, if $A$ is the C-Z constant of $P$ from property $(c)$ of [MPr, Lemma 2.1], integration over annuli and using again that $\overline{y}\not\in H_{\pazocal{D}}$ yield

### v4

@@ 3.3. An equivalent variational capacity
$\gamma_{\text{{sy}},\psi_{\tau},+}(E)$ $\approx\gamma_{\text{{sy}},\psi_{\tau},+}^{\star}(E)\approx\gamma_{\text{{sy}},\psi_{\tau},2}(E)\approx\gamma_{\text{{sy}},\psi_{\tau},\text{{op}}}(E)$
as well as $\limsup_{\tau\to 0}\gamma_{\text{{sy}},\psi_{\tau},+}(E)\lesssim\gamma_{\text{{sy}},+}(E)\lesssim\liminf_{\tau\to 0}\gamma_{\text{{sy}},\psi_{\tau},+}(E)$. The latter relations for the capacities depending on $\psi_{\tau}$ also hold changing each $\gamma_{\text{{sy}}}$ for $\widetilde{\gamma}$.
@@ 4. The construction of the cubes
Let $H$ be a positive Borel measure supported on $E$ such that for any $\lambda_{0}>0$ and any $\lambda\in[0,\lambda_{0}]$, $\mu_{\lambda}:=\mu_{0}+\lambda H\in\Sigma_{n}(E)$. Then,
@@
$\hskip-28.45274pt\geq\frac{\mu_{0}(E)^{2}+2\lambda\mu_{0}(E)H(E)+\lambda^{2}H(E)^{2}}{\mu_{0}(E)+\lambda H(E)+\int_{E}(\pazocal{S}\mu_{0})^{2}\mathop{}\!\mathrm{d}\mu_{0}+\lambda\big{[}\cdots\big{]}+\lambda^{2}\big{[}\cdots\big{]}+\lambda^{3}\big{[}\cdots\big{]}},$
@@
Assume that $\lambda>0$ and divide both sides of the previous inequality by this factor and make $\lambda\to 0$ to obtain the desired estimate. Let us notice that the kernel associated to $\pazocal{S}$ is continuous, meaning that the terms in $\cdots$] and $\int_{E}(\pazocal{S}H)^{2}\mathop{}\!\mathrm{d}H$ are finite, since $H(E)<\infty$ due to the fact that $\mu_{\lambda}\in\Sigma_{n}(E)$ for any $\lambda\in[0,\lambda_{0}]$ with $\lambda_{0}>0$. ∎
@@
We will now prove that for all sufficiently small $\lambda>0$, $\mu_{0}+\lambda\pazocal{L}^{n+1}|_{G}$ belongs to $\Sigma_{n}(E)$, i.e.
@@
where for the last inequality we have applied the reasoning of the second case. Hence, in general, we deduce that for $\lambda>0$ small enough, $\mu_{\lambda}\in\Sigma_{n}(E)$.
@@ 4.1. The construction
since $P_{\text{{sy}}}$ and the measure $\mu$ are nonnegative. Therefore, $\pazocal{S}_{\mu}$ becomes a CZ convolution operator (with continuous kernel), so that it satisfies the weak estimate
@@
Therefore, using the CZ estimates for $P_{\text{{sy}}}\psi_{\tau_{0}}$, the nonnegativity of the latter kernel and relations (4.3) we have
@@ 1. $\text{{{A}}}_{\boldsymbol{1}}$:
    $E$ is contained in the unit ball and consists of a finite union of closed dyadic cubes belonging to a dyadic grid in $\mathbb{R}^{n+1}$ ((with sides parallel to the coordinate axes)\,), all with the same size and with disjoint interiors.
Let us verify that if we deduce the comparability between $\widetilde{\gamma}$ and $\widetilde{\gamma}_{+}$ for $E$ satisfying $\text{{{A}}}_{\boldsymbol{2}}$, we obtain the same result for a general $E$. So fix $E\subset\mathbb{R}^{n+1}$ any compact set and let $\overline{\pazocal{U}_{\delta}(E)}$ be the closed $\delta$-neighborhood of $E$. Consider a grid of dyadic cubes in $\mathbb{R}^{n+1}$ with sides parallel to the axes and diameter smaller than $\delta/2$. Name $E_{0}$ the collection of dyadic cubes that intersect $E$ and notice that $E\subset E_{0}\subset\overline{E_{0}}\subset\overline{\pazocal{U}_{\delta}(E)}$. Now we would have
$\widetilde{\gamma}(E)\leq\widetilde{\gamma}(\overline{E_{0}})\lesssim\widetilde{\gamma}_{+}(\overline{E_{0}})\leq\widetilde{\gamma}_{+}(\overline{\pazocal{U}_{\delta}(E)})\leq\gamma_{+}(\overline{\pazocal{U}_{\delta}(E)}).$
@@
Finally, we present a last assumption that motivates the title of this section:
@@
The aim of this section is to prove the comparability between $\widetilde{\gamma}$ and $\widetilde{\gamma}_{+}$ under this last assumption $\text{{{A}}}_{\boldsymbol{3}}$.
@@ 5.1. Basic definitions and properties
$\bigg{\rvert}\int_{Q}\varphi(\overline{x})\mathop{}\!\mathrm{d}\overline{x}\bigg{\rvert}\leq\bigg{\rvert}\int_{Q}\partial_{t}(P\ast\varphi)(\overline{x})\mathop{}\!\mathrm{d}\overline{x}\bigg{\rvert}+\bigg{\rvert}\int_{Q}(\Delta_{x})^{1/2}(P\ast\varphi)(\overline{x})\mathop{}\!\mathrm{d}\overline{x}\bigg{\rvert}$
@@
$\bigg{\rvert}\int_{Q}\partial_{t}(P\ast\varphi)(\overline{x})\mathop{}\!\mathrm{d}\overline{x}\bigg{\rvert}$ $=\bigg{\rvert}\int_{Q_{1}}\Big{(}P\ast\varphi(x,b)-P\ast\varphi(x,a)\Big{)}\mathop{}\!\mathrm{d}x\bigg{\rvert}\lesssim\|P\ast\varphi\|_{\infty}\pazocal{L}^{n}(Q_{1})\leq\ell(Q)^{n}.$
@@
$(-\Delta)^{1/2}\approx\sum_{j=1}^{n}\partial_{j}R_{j}$
where $R_{j},\,1\leq j\leq n$ are the Riesz transforms with Fourier multiplier $\xi_{j}/|\xi_{j}|$. Notice that if we prove that $P\ast(R_{j}\varphi)\in L^{\infty}(\mathbb{R}^{n+1})$ for each $j=1,\ldots,n$, we are done, since we may argue as above decomposing the cube $Q$ as a cartesian product of an interval and an $n$-dimensional cube. To study $P\ast(R_{j}\varphi)$ we consider its Fourier transform $\frac{\xi_{j}}{|\xi|}\widehat{P}\widehat{\varphi}$. Notice that $P\in(L^{1}+L^{2})(\mathbb{R}^{n+1})$ (just decompose it as $P\chi_{|\overline{x}|\leq 1}+P\chi_{|\overline{x}|>1}$), and therefore $\widehat{P}\in(L^{\infty}+L^{2})(\mathbb{R}^{n+1})$, and we write its decomposition as $\widehat{P}_{\infty}+\widehat{P}_{2}$. Now it is immediate to check that $\frac{\xi_{j}}{|\xi|}\widehat{P}_{\infty}\widehat{\varphi}\in L^{1}(\mathbb{R}^{n+1})$ and we also have $\frac{\xi_{j}}{|\xi|}\widehat{P}_{2}\widehat{\varphi}\in L^{1}(\mathbb{R}^{n+1})$ using Cauchy-Schwarz inequality. Therefore the Fourier transform of $P\ast(R_{j}\varphi)$ is integrable, implying $P\ast(R_{j}\varphi)\in L^{\infty}(\mathbb{R}^{n+1})$ that is what we wanted to see.
@@ 5.2. The exceptional sets $\boldsymbol{H_{\pazocal{D}}}$ and $\boldsymbol{T_{\pazocal{D}}}$
Pick a subfamily of disjoint maximal cubes $\{Q_{k}\}_{k}$ from $\pazocal{D}_{H}$ so that we may define the exceptional set HDH{D} as follows
@@
We define the exceptional set TDT{D} in a similar manner but imposing a proper accretivity condition on the Radon-Nikodym derivative $b$. Consider the family $\pazocal{D}_{T}\subset\pazocal{D}$ of cubes satisfying
@@
5.3. Verifying the hypotheses of a TbTb\,-theorem
The goal of this subsection is to prove the lemma below, analogous to [To3, Lemma 6.8 (Main Lemma)]. We remark that in order to apply the $Tb$-theorem found in [NTVo1] we will still need to check an additional weak boundedness property for a particular suppressed kernel, due to the lack of anti-symmetry of our kernel $P$. In the proof of the aforementioned reference or that of [To3, Theorem 5.1], such condition needs not to be verified, by the anti-symmetry of the Cauchy and Riesz kernels.
@@ 5.3. Verifying the hypotheses of a TbTb\,-theorem
    If $\mu(B_{r}(\overline{x}))>Lr^{n}$ ((\,for some big constant $L)$, then $B_{r}(\overline{x})\subset H$. In particular, for any $\overline{x}\in F\setminus{H}$ and $r>0$, one has $\mu(B_{r}(\overline{x}))\leq Lr^{n}$.
@@
Many of the hypothesis in Lemma 5.3 are verified for our particular choice of $\mu$ and $\nu$. Properties 1, 2 and 4 are proved in Lemma 5.1, and 3 follows from the third property of the same lemma, and by assumption $\text{{{A}}}_{\boldsymbol{3}}$ combined with $r_{i}^{n}\approx\widetilde{\gamma}(2Q_{i}\cap E)$. Moreover, taking $H\subset F$ to be as in (5.6), properties 6 and 7 are also guaranteed by definition. So it remains to check hypothesis 5. To do so, let us first present a series of auxiliary results.
@@
To prove 2, consider $\psi$ as before and define
@@
that is such that $\|\nabla\eta\|_{\infty}\leq\ell(2Q)^{-1}$. By Remark 2.1, we shall apply [MPr, Theorem 3.1] to deduce $\|P\ast\eta T\|_{\infty}\lesssim 1$ and $\|P^{\ast}\ast\eta T\|_{\infty}\lesssim 1$. Hence $|\langle\eta T,1\rangle|\lesssim\widetilde{\gamma}(E)$, meaning
@@
$|P\ast\widetilde{\varphi_{i}T_{0}}|(\overline{x})\leq\frac{|\langle T_{0},\varphi_{i}\rangle|}{\pazocal{L}^{n+1}(B_{i})}\int_{B_{i}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}\lesssim\frac{1}{r_{i}}\int_{B_{i}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}},$
@@
$\int_{B_{i}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}$ $=\int_{B_{i}\cap\{|\overline{x}-\overline{y}|\geq r_{i}\}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}+\int_{B_{i}\cap\{|\overline{x}-\overline{y}|<r_{i}\}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}$
$\leq\frac{\pazocal{L}^{n+1}(B_{i})}{r_{i}^{n}}+\int_{B_{4r_{i}}(\overline{x})}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}\lesssim r_{i}.$
@@
(5.10) $\hskip 56.9055pt+\frac{|\langle\varphi_{i}T_{0},1\rangle|}{\pazocal{L}^{n+1}(B_{i})}\int_{B_{i}}\big{\rvert}P(\overline{z}-\overline{y})-P(\overline{z}-c_{B_{i}})\big{\rvert}\mathop{}\!\mathrm{d}\overline{y}.$
@@
$\frac{|\langle\varphi_{i}T_{0},1\rangle|}{\pazocal{L}^{n+1}(B_{i})}\int_{B_{i}}\big{\rvert}P(\overline{z}-\overline{y})$- $P(\overline{z}-c_{B_{i}})\big{\rvert}\mathop{}\!\mathrm{d}\overline{y}$
$\lesssim\frac{1}{r_{i}}\frac{1}{|\overline{z}-c_{B_{i}}|^{n+1}}\int_{B_{i}}|\overline{y}-c_{B_{i}}|\mathop{}\!\mathrm{d}\overline{y}$
@@
[…]
@@
[…]
@@
$|R_{\varepsilon}(\overline{x}_{0})|\leq\frac{1}{\varepsilon^{n+1}}\int_{B_{\varepsilon}(0)}\frac{\mathop{}\!\mathrm{d}\overline{x}}{|\overline{x}_{0}-\overline{x}|^{n}}\leq\frac{1}{\varepsilon^{n+1}}\int_{2B_{\varepsilon}(\overline{x}_{0})}\frac{\mathop{}\!\mathrm{d}\overline{x}}{|\overline{x}_{0}-\overline{x}|^{n}}\lesssim\frac{1}{\varepsilon^{n}},$
@@
$|R_{\varepsilon}\ast\alpha_{i}(\overline{z})|\leq\int_{\text{supp}(\Psi_{\varepsilon}\ast\alpha_{i})}P(\overline{z}-\overline{y})\big{\rvert}\Psi_{\varepsilon}\ast\alpha_{i}(\overline{y})\big{\rvert}\mathop{}\!\mathrm{d}\overline{y}.$
@@
$|R_{\varepsilon}\ast\alpha_{i}(\overline{z})|$ $\leq\|\Psi_{\varepsilon}\ast\alpha_{i}\|_{\infty}\int_{\pazocal{U}_{\varepsilon}(2Q_{i})}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{z}-\overline{y}|^{n}}\lesssim(\ell(Q_{i})+\varepsilon)\cdot\|\Psi_{\varepsilon}\ast\alpha_{i}\|_{\infty}$
@@
$\hskip-5.69046pt=\sum_{i=1}^{N}\bigg{(}\int_{4Q_{i}}R_{\ast}\ast\big{(}\widetilde{\varphi_{i}T_{0}}-\varphi_{i}T_{0}\big{)}(\overline{y})\mathop{}\!\mathrm{d}\mu(\overline{y})+\int_{F\setminus{(4Q_{i}\cup H)}}R_{\ast}\ast\big{(}\widetilde{\varphi_{i}T_{0}}-\varphi_{i}T_{0}\big{)}(\overline{y})\mathop{}\!\mathrm{d}\mu(\overline{y})\bigg{)}$
$\hskip-5.69046pt=:\sum_{i=1}^{N}I_{i,1}+I_{i,2},$
@@
The proof of 5 concludes that of Lemma 5.3. In order to apply the $Tb$-theorem of [NTVo1] or [To3, Theorem 5.1] we still need to check an additional weak boundedness property for our $n$-dimensional CZ kernel, which fails being anti-symmetric.
@@
$\pazocal{S}_{1}^{\prime}$ $:=\big{\{}\overline{x}\in F\,:\,\pazocal{P}_{\ast}\nu(\overline{x})>\alpha\big{\}},$
$\pazocal{S}_{2}^{\prime}$ $:=\big{\{}\overline{x}\in F\,:\,\pazocal{P}^{\ast}_{\nu,\ast}1(\overline{x})>\alpha\big{\}},$
@@
[…]
@@
Regarding the third term, if $A$ is the CZ constant of $P$ from property $(c)$ of [MPr, Lemma 2.1], integration over annuli and using again that $\overline{y}\not\in H_{\pazocal{D}}$ yield